### v2

# Bijective rigidity of uniform Roe algebras and injectivity of the comparison map

Abstract.
We show that, for uniformly locally finite metric spaces $X$ and $Y$ with isomorphic uniform Roe algebras $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$, the existence of a bijective coarse equivalence $f\colon X\to Y$ is equivalent to the injectivity of the 0th comparison map appearing in the HK conjecture for coarse groupoids.
We further prove that the 0th comparison map is injective unconditionally. Moreover, if the underlying space is coarsely connected, this map is in fact split-injective.
Introduction
Coarse geometry studies metric spaces by focusing on their large-scale structure while disregarding small-scale features. Since local properties play no essential role in this setting, it is customary to restrict attention to discrete metric spaces. Let $(X,d)$ and $(Y,\partial)$ be metric spaces. A map $f\colon X\to Y$ is called coarse if for every $R\geq 0$ there exists $S\geq 0$ such that $f$ maps subsets of $X$ of diameter at most $R$ to subsets of $Y$ of diameter at most $S$. Coarse maps are precisely those maps that preserve the large-scale geometry of metric spaces. Two coarse maps $f,g\colon X\to Y$ are said to be close if they are at bounded distance, that is,
$\sup_{x\in X}\partial\bigl(f(x),g(x)\bigr)<\infty.$
Throughout this paper, we work with uniformly locally finite metric spaces (see Definition 1.1), also referred to as bounded geometry metric spaces. A classical example is provided by a Cayley graph of a finitely generated group with the shortest path metric. The coarse category $\mathbf{Coarse}$ is the category whose objects are uniformly locally finite metric spaces and whose morphisms are closeness classes of uniformly finite-to-one [^1] A map $f\colon X\to Y$ is uniformly finite-to-one if there exists $N\geq 1$ such that the preimage of every point of $Y$ consists of at most $N$ elements (see Section 1.1). Coarse equivalences are automatically uniformly finite-to-one. coarse maps. Isomorphisms in $\mathbf{Coarse}$ are called coarse equivalences. In the case of finitely generated groups equipped with word metrics, the notion of coarse equivalence coincides with quasi-isometry. We say that two uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ are (bijectively) coarsely equivalent if there exists a (bijective) coarse equivalence between them.
Coarse geometry is intrinsically connected to the theory of étale groupoids [^2] We refer the reader to [22] for a detailed introduction to étale groupoids and their $C^{*}$-algebras. . An étale groupoid is a topological groupoid $\mathcal{G}$ with unit space $\mathcal{G}^{(0)}$ such that the source and range maps $s,r\colon\mathcal{G}\to\mathcal{G}^{(0)}$ are local homeomorphisms. An étale groupoid $\mathcal{G}$ is said to be ample if its unit space $\mathcal{G}^{(0)}$ is totally disconnected. The theory of étale groupoids has attracted significant attention due to its deep connections with topological dynamics, number theory, and, in particular, operator algebras. To any étale groupoid $\mathcal{G}$ one can associate a $C^{*}$-algebra $C^{*}_{r}(\mathcal{G})$, called the reduced groupoid $C^{*}$-algebra of $\mathcal{G}$. In [14], Li showed that every classifiable [^3] By a “classifiable $C^{*}$-algebra” we mean a unital, separable $C^{*}$-algebra with finite nuclear dimension that satisfies UCT. simple $C^{*}$-algebra arises as the reduced groupoid $C^{*}$-algebra of a topologically principal (possibly twisted) étale groupoid. In [8], Crainic and Moerdijk introduced a homology theory for étale groupoids, which is now a substantial tool for analysing the $K$-theory of the reduced groupoid $C^{*}$-algebras arising from ample groupoids. Precisely, in [17] Matui showed that, for groupoids arising from shifts of finite type, the $K_{0}$-groups (respectively, $K_{1}$-groups) of the associated reduced groupoid $C^{*}$-algebras are isomorphic to the direct sum of the even (respectively, odd) homology groups. He conjectured that this phenomenon extends to more general ample groupoids.
Conjecture (HK conjecture).
Let $\mathcal{G}$ be a minimal, essentially principal, ample groupoid with compact unit space. For =0,1*=0,1, there is an isomorphism
$K_{*}\bigl(C^{*}_{r}(\mathcal{G})\bigr)\cong\bigoplus_{i=0}^{\infty}H_{2i+*}(\mathcal{G};\mathbb{Z}),$
where $H_{*}(\mathcal{G};\mathbb{Z})$ are the Crainic and Moerdijk homology groups of $\mathcal{G}$.
The HK conjecture has been verified for many classes of ample groupoids, in some cases even beyond the stated hypotheses. However, counterexamples to the HK conjecture were constructed by Scarparo [21] and Deeley [9]. It is well known that for any ample groupoid $\mathcal{G}$ and $i=0,1$ there exists canonical group homomorphism
$\alpha_{i}\colon H_{i}(\mathcal{G};\mathbb{Z})\longrightarrow K_{i}\bigl(C^{*}_{r}(\mathcal{G})\bigr),$
called the $i$th comparison map (see Definition 1.15 for the definition of $\alpha_{0}$ in the case of coarse groupoids). At present, only partial results are known concerning the existence and properties of comparison maps in higher degrees.
For a uniformly locally finite metric space $(X,d)$, the authors of [23] constructed a Hausdorff, principal, ample groupoid $\mathcal{G}(X,d)$, called the coarse groupoid of $(X,d)$, whose unit space is the Stone-Cech compactification of $X$. The construction of the coarse groupoid yields a functor from the coarse category to the category of étale groupoids with generalised groupoid homomorphisms. The authors of [5] proved that the HK-conjecture holds for coarse groupoids associated with uniformly locally finite metric spaces of sufficiently small asymptotic dimension. Moreover, they showed that the groupoid homology of the coarse groupoid $\mathcal{G}(X,d)$ is canonically isomorphic to the uniformly finite homology (see Definition 1.10) of $(X,d)$, which will be denoted by $H_{i}^{\mathrm{uf}}(X;\mathbb{Z})$. The reduced groupoid $C^{*}$-algebra of a coarse groupoid associated to a uniformly locally finite metric space $(X,d)$ is more commonly known as the uniform Roe algebra $C^{*}_{u}(X)$. The latter has a simple description in terms of bounded operators on the Hilbert space $\ell^{2}(X)$. A bounded operator $T\in\mathcal{B}(\ell^{2}(X))$ is said to have controlled propagation if the quantity
$\operatorname{prop}(T)=\sup\bigl\{d(x,y)\,\bigm|\,(x,y)\in X\times X\text{ such that }\langle T\delta_{x},\delta_{y}\rangle\neq 0\bigr\}$
is finite. The uniform Roe algebra $C^{*}_{u}(X)$ is defined as the norm closure in $\mathcal{B}(\ell^{2}(X))$ of the *-algebra of bounded operators with controlled propagation. If $G$ is a finitely generated discrete group equipped with a word metric, then the uniform Roe algebra of $G$ is canonically isomorphic to the reduced crossed product $\ell^{\infty}(G)\rtimes_{r}G$. We refer to [20, Proposition 10.29] for an explicit construction of the isomorphism between $C^{*}_{u}(X)$ and $C^{*}_{r}(\mathcal{G}(X,d))$.
Uniform Roe algebras encode a substantial amount of coarse-geometric information about the underlying space. For example, a uniformly locally finite metric space $(X,d)$ has property A if and only if $C^{*}_{u}(X)$ is a nuclear $C^{*}$-algebra [23], and $(X,d)$ is amenable (as a metric space) if and only if $C^{*}_{u}(X)$ admits a tracial state [1]. This naturally leads to the question of whether the uniform Roe algebra fully determines the coarse geometry of the underlying space. The authors of [2] showed that the Morita equivalence class of the uniform Roe algebra is a complete invariant of coarse equivalence.
Theorem 0.1 (Baudier, Braga, Farah, Khukhro, Vignati, Willett, 2021).
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. The $C^{*}$-algebras $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$ are Morita equivalent if and only if $(X,d)$ and $(Y,\partial)$ are coarsely equivalent.
An analogous statement for isomorphisms of uniform Roe algebras and bijective coarse equivalences, known as “the bijective rigidity problem”, was a long-standing conjecture in coarse geometry. It was recently resolved by Vignati in [27].
Theorem 0.2 (Vignati, 2026).
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. The $C^{*}$-algebras $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$ are isomorphic if and only if $(X,d)$ and $(Y,\partial)$ are bijectively coarsely equivalent.
The $K$-theory of the uniform Roe algebra is of particular importance both from a mathematical perspective and for applications in physics. In [25], Špakula introduced the uniform coarse Baum–Connes conjecture, which asks whether a certain assembly map from the uniform $K$-homology of a metric space to the $K$-theory of the associated uniform Roe algebra is an isomorphism. On the other hand, the Baum–Connes conjecture for ample groupoids yields another assembly map, from the topological $K$-theory of the coarse groupoid to the $K$-theory of the associated uniform Roe algebra, which was conjectured [^4] Counterexamples to surjectivity were constructed in [12]. to be an isomorphism. It was shown in [11] that, when the uniformly locally finite metric space arises from a finitely generated group, the two conjectures are equivalent. In [13], Kubota used uniform Roe algebras as models for controlled topological phases of bulk and edge quantum systems. In particular, the $K$-theory of the uniform Roe algebra plays a decisive role in the study of the so-called bulk index.
Main results
In this paper, we establish a connection between the bijective rigidity problem and the injectivity of the 0th comparison map appearing in the HK conjecture for coarse groupoids.
We say that a collection $\mathcal{C}$ of metric spaces is admissible (see Definition 3.1) if it is stable under taking subspaces and doublings [^5] On the level of coarse groupoids, a doubling corresponds to taking the product with $\mathcal{R}_{n}$, where $\mathcal{R}_{n}$ denotes the discrete groupoid associated to the full equivalence relation on $n$ points. See Section 1.1 for the coarse geometric definition. . For a property $\mathrm{P}$ of metric spaces, we say that an admissible collection $\mathcal{C}$ satisfies $\mathrm{P}$ if every element of $\mathcal{C}$ satisfies $\mathrm{P}$. For example, since property A passes to subspaces and doublings, the collection of metric spaces satisfying property A forms an admissible collection. The collection of all uniformly locally finite metric spaces, clearly, forms an admissible collection. The following lemma constitutes the main result of this paper.
Lemma A.
Let $\mathcal{C}$ be an admissible collection of uniformly locally finite metric spaces. The following statements are equivalent:
  1. (1)
    For every uniformly locally finite metric space $(X,d)$ in $\mathcal{C}$, the 0th comparison map $\alpha_{0}\colon H^{\mathrm{uf}}_{0}(X;\mathbb{Z})\to K_{0}\bigl(C^{*}_{u}(X)\bigr)$ is injective;
  2. (2)
    For any uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ such that $C^{*}_{u}(Y)\cong C^{*}_{u}(X)$, the coarse equivalence given by Theorem 0.1 is close to a bijective coarse equivalence.
By Theorem 0.2, the collection of all uniformly locally finite metric spaces satisfies the second condition of Lemma A; therefore, the first assertion holds unconditionally. Nevertheless, we formulate Lemma A as establishing a relationship between the two conjectural statements, as this viewpoint clarifies the techniques underlying our arguments. Moreover, we expect that the methods developed here can be adapted to broader classes of ample groupoids, potentially yielding new results on the injectivity of comparison maps and on rigidity phenomena for ample groupoids.
Theorem B.
Let $(X,d)$ be a uniformly locally finite metric space, and $\mathcal{G}$ be its coarse groupoid. The comparison map $\alpha_{0}\colon H_{0}(\mathcal{G};\mathbb{Z})\to K_{0}(C^{*}_{r}(\mathcal{G}))$ is injective. Moreover, if $(X,d)$ is coarsely connected, then $\alpha_{0}$ is split injective.
The above theorem partially extends the results of [5], where the authors proved that $\alpha_{0}$ is split injective for uniformly locally finite metric spaces of small asymptotic dimension. To the best of our knowledge, no further results concerning the (split) injectivity of $\alpha_{0}$ for uniform Roe algebras are currently available. There exist many examples of uniformly locally finite metric spaces with infinite asymptotic dimension that are not coarsely connected, such as expander graphs. By contrast, any finitely generated group equipped with a word-length metric is coarsely connected. In this setting, for a finitely generated group $G$, the uniformly finite homology of $G$ is canonically isomorphic to the group homology of $G$ with coefficients in $\ell^{\infty}(G,\mathbb{R})$ (see [10, Proposition A.10]).
Corollary C.
Let $G$ be a finitely generated group equipped with the word-length metric. The canonical map
$\alpha_{0}^{G}\colon H_{0}\bigl(G;\ell^{\infty}(G,\mathbb{R})\bigr)\longrightarrow K_{0}\bigl(\ell^{\infty}(G)\rtimes_{r}G\bigr)$
is split-injective.
Recent results of [19] show that the rational HK conjecture holds for all second-countable, locally compact, Hausdorff, ample groupoids that satisfy the rational Baum–Connes conjecture, and have torsion-free stabilisers. Coarse groupoids fall outside this framework: they are not second-countable and do not in general satisfy the Baum–Connes conjecture (see [12]). Nevertheless, by tensoring with $\mathbb{Q}$ in Theorem B, we immediately obtain unconditional split injectivity of the rational comparison map (see Remark 3.6).
Outline of the paper
In Section 1, we introduce the necessary background from coarse geometry, uniform Roe algebras, and uniformly finite homology. We briefly recall the construction of the coarse groupoid from [23], sketch the main ideas underlying the proof of Theorem 0.1, review the results on uniformly finite homology, and construct the 0th comparison map.
Section 2 develops a framework of covering isometries for uniform Roe algebras, analogous to the construction for Roe algebras (see [30, Section 4.3]). Let $H$ be a separable Hilbert space. A uniform cover of a uniformly finite-to-one coarse map $f\colon X\to Y$ is an isometry
$V_{f}\colon\ell^{2}(X,H)\longrightarrow\ell^{2}(Y,H)$
which acts fibrewise in a manner compatible with the action of $f$ on the underlying metric spaces, and is well-behaved with respect to finite-dimensional subspaces of the Hilbert space $H$ (see conditions $(2)$ and $(3)$ of Definition 2.1). This notion already appears implicitly in [7] in the special case of coarse equivalences. We show that any uniformly finite-to-one coarse map $f\colon(X,d)\to(Y,\partial)$ admits a uniform cover, and that the map $\operatorname{Ad}_{V_{f}}$ is a *-homomorphism between the stabilisations of the corresponding uniform Roe algebras. Moreover, the induced map in $K$-theory depends only on the closeness class of $f$. We further show that the 0th comparison map defines a natural transformation from uniformly finite homology to the $K_{0}$-theory of uniform Roe algebras. The section concludes with a brief discussion of this framework from the perspective of étale groupoids.
In Section 3, we prove Lemma A. The key step in the proof of the implication $(1)\Rightarrow(2)$ is to show that, given an isomorphism of uniform Roe algebras $\Phi\colon C^{*}_{u}(X)\longrightarrow C^{*}_{u}(Y)$, the induced map in $K$-theory coincides with the map arising from the coarse equivalence $f_{\Phi}\colon X\longrightarrow Y$ provided by Theorem 0.1. Since $K_{0}(\Phi)$ preserves the class of the unit, it follows that
$H_{0}^{\mathrm{uf}}(f_{\Phi})([\mathbbm{1}_{X}])=[\mathbbm{1}_{Y}]$
modulo $\ker(\alpha_{0}^{Y})$. As the comparison maps are assumed to be injective, the result of [29, Theorem A] (see Theorem 1.12) completes the proof of the implication $(1)\Rightarrow(2)$. For the converse implication, to each positive function $h\in\ell^{\infty}(X,\mathbb{Z})$ we associate a subspace $X(h)$ of a doubling of $X$. We show that if $h_{1},h_{2}$ are positive functions such that
$\alpha_{0}^{X}([h_{1}])=\alpha_{0}^{X}([h_{2}]),$
then the uniform Roe algebras of $X(h_{1})$ and $X(h_{2})$ are isomorphic. By the already established implication, there exists a bijective coarse equivalence $f\colon X(h_{1})\to X(h_{2})$. The graph of $f$ provides a cycle $\delta\in C_{1}^{\mathrm{uf}}(X;\mathbb{Z})$ such that
$\partial_{1}(\delta)=[h_{1}]-[h_{2}].$
This completes the proof of Lemma A. In the final section, we prove Theorem B and discuss further consequences of Lemma A.
Acknowledgements
The author is grateful to his advisors, Alessandro Vignati and Romain Tessera, for their advice on the presentation of this work. He also thanks Ján Špakula and the University of Southampton for their hospitality during his research visit. The author is indebted to Rufus Willett for valuable comments on Remark 4.3. This work was partially supported by the travel grant ECOST Actions CaLISTA (E-COST-GRANT-CA21109-15f1764d) and by the ANR grant ROAR (ANR-25-CE40-5029) .
1. Preliminaries
In this section, we introduce the basic notions from coarse geometry, uniform Roe algebras, and uniformly finite homology. For a more detailed exposition, we refer the reader to [20]. Let $X$, $Y$, and $Z$ be sets. For a subset $E\subset X\times Y$ we define its adjoint as
$E^{-1}:=\{(y,x)\mid(x,y)\in E\}.$
For subsets $E\subset X\times Y$ and $F\subset Y\times Z$ we define their composition as
$E\circ F:=\{(x,z)\mid\text{for some }y\in Y\text{ such that }(x,y)\in E,\,(y,z)\in F\}.$
The subset $\Delta_{X}=\{(x,x)\mid x\in X\}$ of $X\times X$ is called the diagonal of $X$. Given a map $f\colon X\to Y$, we regard its graph $\operatorname{Graph}(f)$ as a subset of $Y\times X$. By a partially defined map $t\colon X\dashrightarrow Y$ we mean a triple
$\bigl(\operatorname{dom}(t),t,\operatorname{ran}(t)\bigr),$
where $\operatorname{dom}(t)$ and $\operatorname{ran}(t)$ are subsets of $X$ and $Y$, respectively, and $t$ is map from $\operatorname{dom}(t)$ to $\operatorname{ran}(t)$.
1.1. Coarse geometry
Coarse geometry studies geometric objects from the perspective of their large-scale structure. One of its principal objects of interest is the class of uniformly locally finite metric spaces.
Definition 1.1.
A metric space $(X,d)$ is said to be uniformly locally finite if for every $R\geq 0$ there exists $N\in\mathbb{N}$ such that
$|B_{R}(x)|\leq N$
for all $x\in X$, where $B_{R}(x)$ denotes the ball of radious $R$ centered at $x$.
In the literature, uniformly locally finite metric spaces are also referred to as bounded geometry spaces. For example, any connected uniformly locally finite graph is a uniformly locally finite metric space, when equipped with the shortest path metric; in particular, the Cayley graph of a finitely generated group provides a canonical example. Discretisations of bounded geometry Riemannian manifolds yield further examples. To define appropriate morphisms between uniformly locally finite metric spaces, the notion of asymptotic sets plays a central role.
Definition 1.2.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. Two subsets $S_{1},S_{2}\subset Y\times X$ are said to be asymptotic, denoted $S_{1}\asymp S_{2}$, if there exists $R\geq 0$ such that
$S_{2}\subset\bigcup_{(y,x)\in S_{1}}\bigl(B_{R}(y)\times B_{R}(x)\bigr)\quad\text{and}\quad S_{1}\subset\bigcup_{(y,x)\in S_{2}}\bigl(B_{R}(y)\times B_{R}(x)\bigr).$
Informally, two subsets of $Y\times X$ are asymptotic if each is contained in a bounded neighbourhood of the other. A subset of $X\times X$ that is asymptotic to the diagonal $\Delta_{X}$ is called an entourage. It is straightforward to verify that asymptoticity defines an equivalence relation. Moreover, for metric spaces $(X,d)$ and $(Y,\partial)$ and maps $f,g\colon X\to Y$, the graphs $\operatorname{Graph}(f)$ and $\operatorname{Graph}(g)$ are asymptotic if and only if the maps are uniformly close, that is,
$\sup_{x\in X}\partial\bigl(f(x),g(x)\bigr)<\infty.$
In this case, we say that $f$ is close to $g$. The notion of closeness is compatible with composition: the closeness class of a composition depends only on the closeness classes of the constituent maps.
Definition 1.3.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. A map $f\colon X\to Y$ is said to be coarse if for every $R\geq 0$ there exists $S\geq 0$ such that
$f\bigl(B_{R}(x)\bigr)\subset B_{S}\bigl(f(x)\bigr)\quad\text{for all }x\in X.$
A coarse map $f\colon X\to Y$ is called a coarse equivalence if there exists a coarse map $g\colon Y\to X$ such that the compositions $f\circ g$ and $g\circ f$ are close to $\operatorname{id}_{Y}$ and $\operatorname{id}_{X}$, respectively. In this case, $g$ is called a coarse inverse of $f$. A coarse map $f\colon X\to Y$ is said to be uniformly finite-to-one if
$\sup_{y\in Y}\bigl|f^{-1}(\{y\})\bigr|<\infty.$
We say that two uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ are (bijectively) coarsely equivalent if there exists a (bijective) coarse equivalence
$f\colon(X,d)\longrightarrow(Y,\partial).$
It is straightforward to verify that every coarse equivalence between uniformly locally finite metric spaces is uniformly finite-to-one. We define the category $\mathbf{Coarse}$ to have uniformly locally finite metric spaces as objects and closeness classes of uniformly finite-to-one coarse maps as morphisms. Isomorphisms in this category are precisely the coarse equivalences.
Given a uniformly locally finite metric space $(X,d)$ and a subset $A\subset X$, the restriction of the metric $d$ to $A$ turns $(A,d)$ into a uniformly locally finite metric space. Another construction that will play an important role is that of a doubling. For a uniformly locally finite metric space $(X,d)$, its $n$th doubling is the uniformly locally finite metric space $(X^{(n)},d_{n})$ defined by
$X^{(n)}:=X\times\{1,\ldots,n\},\qquad d_{n}\bigl((x,i),(y,j)\bigr):=d(x,y)+|i-j|,$
for $x,y\in X$ and $1\leq i,j\leq n$. It is straightforward to check that, for each $1\leq k\leq n$, the inclusion
$i_{k}\colon X\longrightarrow X^{(n)},\qquad x\longmapsto(x,k),$
is a coarse equivalence. The projection map $p\colon X\times\{1,\ldots,n\}\to X$ constitutes its coarse inverse.
For a uniformly locally finite metric space $(X,d)$, the authors of [23] constructed a topological groupoïd that captures the coarse geometry of $(X,d)$. In this paper, we do not appeal to the theory of topological groupoids as a primary tool in our proofs. Nevertheless, we believe that it is instructive to highlight the connection between coarse geometry and topological groupoids. For a detailed introduction to topological groupoids, we refer the reader to [22]. A groupoid is a small category in which every morphism is invertible. We shall identify a groupoid with its set of morphisms, denoted by $\mathcal{G}$. The set of objects, or units, is denoted by $\mathcal{G}^{(0)}$ and is identified with the subset of identity morphisms in $\mathcal{G}$. The groupoid $\mathcal{G}$ is equipped with the range and source maps
$r,s\colon\mathcal{G}\longrightarrow\mathcal{G}^{(0)},$
which assign to each arrow its range and source, respectively, together with a partially defined multiplication
$\mathcal{G}{{}_{s}}\times_{r}\mathcal{G}:=\{(\gamma_{1},\gamma_{2})\in\mathcal{G}\times\mathcal{G}\mid s(\gamma_{1})=r(\gamma_{2})\}\longrightarrow\mathcal{G},\qquad(\gamma_{1},\gamma_{2})\longmapsto\gamma_{1}\gamma_{2},$
and an inversion map $\mathcal{G}\to\mathcal{G}$, $\gamma\mapsto\gamma^{-1}$. These maps satisfy the usual axioms that make $\mathcal{G}$ into a category. We shall be concerned with topological groupoids. A locally compact, Hausdorff groupoid is a groupoid $\mathcal{G}$ endowed with a locally compact, Hausdorff topology such that the range and source maps, the multiplication, and the inversion are continuous. Here $\mathcal{G}{{}_{s}}\times_{r}\mathcal{G}$ is equipped with the subspace topology inherited from the product topology on $\mathcal{G}\times\mathcal{G}$. A locally compact, Hausdorff groupoid $\mathcal{G}$ is said to be étale [^6] Many authors do not require étale groupoids to be Hausdorff. Throughout this paper, we restrict attention to the Hausdorff case. if its range map (equivalently, its source map) is a local homeomorphism. A groupoid $\mathcal{G}$ is said to be principal if its isotropy bundle
$\operatorname{Iso}(\mathcal{G}):=\{\gamma\in\mathcal{G}\mid s(\gamma)=r(\gamma)\}$
coincides with $\mathcal{G}^{(0)}$, i.e. the only arrows whose source and range coincide are the identity arrows. An étale groupoid is said to be ample if its unit space $\mathcal{G}^{(0)}$ is totally disconnected.
In [23], Skandalis, Tu, and Yu associated to every uniformly locally finite metric space $(X,d)$ a topological groupoid $\mathcal{G}(X,d)$, defined as follows. For a countable discrete topological space $A$, let $\beta A$ denote its Stone–Čech compactification. Given an entourage $E\subset X\times X$, denote by $\overline{E}$ its closure in $\beta(X\times X)$. One then defines a topological space
$\mathcal{G}(X,d)=\bigcup_{\begin{subarray}{c}E\text{ is an}\\ \text{entourage}\end{subarray}}\overline{E}\;\subset\;\beta(X\times X),$
equipped with the weak topology. The pair groupoid $X\times X$ embeds as a dense open subset of $\mathcal{G}(X,d)$. It is shown in [23] that the groupoid operations on $X\times X$ extend continuously to $\mathcal{G}(X,d)$, endowing it with the structure of a Hausdorff, principal, ample groupoid. We shall refer to $\mathcal{G}(X,d)$ as the coarse groupoid associated to $(X,d)$. Note that $\mathcal{G}(X,d)$ is not second countable unless the metric space $(X,d)$ is finite.
1.2. Uniform Roe algebras and rigidity
Uniform Roe algebras are $C^{*}$-algebras associated with uniformly locally finite metric spaces. Let $H$ be a separable Hilbert space and let $X$ be a set. Denote by $\ell^{2}(X,H)$ the Hilbert space of all square-summable functions from $X$ to $H$. There is a canonical identification
$\ell^{2}(X,H)\cong\ell^{2}(X)\otimes H.$
For a subset $A\subseteq X$, let $\mathbbm{1}_{A}$ denote the orthogonal projection onto $\ell^{2}(A,H)$. If $A=\{x\}$ is a singleton, we simply write $\mathbbm{1}_{x}$. The main ingredient in the definition of uniform Roe algebras is the notion of support.
Definition 1.4.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $H$ be a separable Hilbert space. The support of a bounded operator $T\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$ is the set
$\operatorname{supp}(T):=\{(y,x)\in Y\times X\mid\|\mathbbm{1}_{y}T\mathbbm{1}_{x}\|\neq 0\}.$
We say that a bounded operator $T\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$ is controlled if its support is asymptotic to a coarse map. An operator $T\in\mathcal{B}(\ell^{2}(X,H))$ is said to have controlled propagation if its support is asymptotic to $\operatorname{id}_{X}$ (equivalently, if its support is an entourage). It is said to be approximable if it is a norm limit of controlled propagation operators.
It is straightforward to verify that the support behaves well with respect to compositions, adjoints, and sums. More precisely, let $(X,d_{X})$, $(Y,d_{Y})$, and $(Z,d_{Z})$ be uniformly locally finite metric spaces, and let
$T_{1},T_{2}\colon\ell^{2}(X,H)\to\ell^{2}(Y,H),\qquad S\colon\ell^{2}(Y,H)\to\ell^{2}(Z,H)$
be bounded operators. Then the supports of sums, products, and adjoints satisfy the following relations:
$\begin{split}&\operatorname{supp}(T_{1}+T_{2})\subset\operatorname{supp}(T_{1})\cup\operatorname{supp}(T_{2}),\\ &\operatorname{supp}(ST_{1})\subset\operatorname{supp}(S)\circ\operatorname{supp}(T_{1}),\\ &\operatorname{supp}(T_{1}^{*})=\operatorname{supp}(T_{1})^{-1}.\end{split}$
In particular, the controlled propagation operators on $\ell^{2}(X,H)$ form a *-subalgebra of $\mathcal{B}(\ell^{2}(X,H))$. Controlled propagation operators in $\mathcal{B}(\ell^{2}(X,H))$ can be naturally quantified by the extent to which their support is spread. For a bounded operator $T\in\mathcal{B}(\ell^{2}(X,H))$, we define its propagation by
$\operatorname{prop}(T)=\sup\{d(x,x^{\prime})\mid(x,x^{\prime})\in\operatorname{Supp}(T)\}.$
Note that an operator $T$ has controlled propagation if and only if $\operatorname{prop}(T)<\infty$.
Definition 1.5.
Let $(X,d)$ be a uniformly locally finite metric space. The uniform Roe algebra $C^{*}_{u}(X)$ of $(X,d)$ is the norm closure in $\mathcal{B}(\ell^{2}(X))$ of the *-algebra of bounded operators of controlled propagation.
Uniform Roe algebras provide a natural bridge between coarse geometry and operator algebras, as their structural properties encode many coarse invariants. Alternatively, uniform Roe algebras can be defined as the reduced groupoid $C^{*}$-algebras of the coarse groupoids associated to uniformly locally finite metric spaces (see [20, Proposition 10.29]). Note that the uniform Roe algebra of a doubling of $X$ is canonically isomorphic to the matrix algebra over $C^{*}_{u}(X)$, that is, for every $n\geq 1$ one has:
$C^{*}_{u}(X^{(n)})\cong C^{*}_{u}(X)\otimes M_{n}(\mathbb{C}).$
We shall also consider stabilisations of uniform Roe algebras. Let $H$ be a separable, infinite-dimensional Hilbert space. The $C^{*}$-algebra $C^{*}_{u}(X)\otimes\mathbb{K}$ admits a natural faithful representation on $\ell^{2}(X,H)$. For each $x\in X$, consider an isometry
$v_{x}\colon H\longrightarrow\ell^{2}(X,H),\qquad w\longmapsto\delta_{x}\otimes w.$
For a finite-dimensional subspace $W\subset H$ and $R\geq 0$, consider the following set of bounded operators on $\ell^{2}(X,H)$:
$\mathbb{C}[X,R,W]:=\bigl\{T\;\big|\;\operatorname{prop}(T)\leq R\text{ and }v_{x}^{*}Tv_{y}\in\mathcal{B}(W)\text{ for all }x,y\in X\bigr\}.$
The following statement summarises the discussion in [2, Section 4], providing a convenient dense *-subalgebra of $C^{*}_{u}(X)\otimes\mathbb{K}$.
Proposition 1.6 ([2, Section 4]).
Let $(X,d)$ be a uniformly locally finite metric space, and let $H$ be a separable Hilbert space. Then the union of the sets $\mathbb{C}[X,R,W]$ over all finite-dimensional subspaces $W\subset H$ and all $R\geq 0$ is dense in $C^{*}_{u}(X)\otimes\mathbb{K}$.
In particular, any approximable operator $T\in\mathcal{B}(\ell^{2}(X,H))$ for which there exists a finite-dimensional subspace $W\subset H$ such that $v_{x}^{*}Tv_{y}$ belongs to $\mathcal{B}(W)$ for all $x,y\in X$ belongs to the stabilisation $C^{*}_{u}(X)\otimes\mathbb{K}$ of the uniform Roe algebra.
The rigidity problem for uniform Roe algebras asks whether Morita equivalences (respectively, isomorphisms) of uniform Roe algebras induce coarse equivalences (respectively, bijective coarse equivalences) between the underlying uniformly locally finite metric spaces. In [2], the authors provided an affirmative answer to the rigidity problem in the case of Morita equivalences.
Theorem 1.7 ([2, Theorem 1.4]).
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces. Then the $C^{*}$-algebras $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$ are Morita equivalent if and only if $(X,d)$ and $(Y,\partial)$ are coarsely equivalent.
The authors of [16] established an analogue of Theorem 1.7 for a wide class of Roe-type algebras. Their methods are also applicable to stabilisations of uniform Roe algebras. In what follows, we recall the strategy of the proof of Theorem 1.7 from [16] in the special case of an isomorphism $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$.
  1. (1)
    One proves that the isomorphism $\Phi$ is spatially implemented, in the sense that there exists a unitary operator $U\colon\ell^{2}(X)\to\ell^{2}(Y)$ such that $\Phi=\operatorname{Ad}_{U}$ (see [24, Lemma 3.1]);
  2. (2)
    One shows that the unitaries $U$ and $U^{*}$ implementing $\Phi$ are coarse-like [^7] In [16, 15], this notion is also referred to as weakly approximately controlled. (see [6, Theorem 3.5]). Recall from [6, Definition 3.1] that an operator
    $Q\colon\ell^{2}(X)\longrightarrow\ell^{2}(Y)$
    is said to be coarse-like if, for every $R\geq 0$ and every $\varepsilon>0$, there exists $S\geq 0$ such that for every contraction $T\in\mathcal{B}(\ell^{2}(X))$ of propagation less than $R$ the operator $\operatorname{Ad}_{Q}(T)$ is $\varepsilon$-close (in norm) to an operator on $\ell^{2}(Y)$ of propagation at most $S$.
  3. (3)
    One shows that, for every $\delta\in(0,1)$, there exist constants $R,S\geq 0$ such that the subset
    $\quad\qquad f_{\delta,R,S}^{U}=\bigcup\left\{B\times A\;\middle|\;\operatorname{diam}(B)\leq R,\ \operatorname{diam}(A)\leq S,\ \bigl\|\mathbbm{1}_{B}\,U\,\mathbbm{1}_{A}\bigr\|>\delta\right\}$
    of $Y\times X$ is asymptotic to the graph of a coarse equivalence $f_{\Phi}$ between $(X,d)$ and $(Y,\partial)$.
The resulting coarse equivalence $f_{\Phi}\colon(X,d)\to(Y,\partial)$ is uniquely determined up to closeness. We shall refer to $f_{\Phi}$ as the coarse equivalence induced by $\Phi$. The following theorem provides a key tool for relating the coarse equivalence $f_{\Phi}$ back to the isomorphism $\Phi$.
Theorem 1.8 ([16, Theorem 4.5]).
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$ be an isomorphism. Let $U$ be a unitary implementing $\Phi$, and let $f_{\Phi}$ be a coarse equivalence induced by $\Phi$. Then, for an infinite-dimensional Hilbert space $H$, the unitary $U\otimes\operatorname{id}_{H}$ is a norm limit of unitaries supported on subsets of $Y\times X$ that are asymptotic to the graph of $f_{\Phi}$.
1.3. Uniformly finite homology
Uniformly finite homology was introduced by Block and Weinberger in [4] to study tilings of non-amenable polyhedra and characteristic numbers of manifolds whose universal covers admit metrics of positive scalar curvature. It was subsequently employed by Whyte in [29] to prove the geometric von Neumann conjecture.
Definition 1.9.
Let $(X,d)$ be a uniformly locally finite metric space, and let $n\geq 1$ be an integer. A function $f\colon X^{n}\to\mathbb{Z}$ is said to have controlled propagation if there exists $R\geq 0$ such that
$f(x_{1},\ldots,x_{n})=0\quad\text{whenever}\quad\max_{i,j}d(x_{i},x_{j})>R.$
Let $(X,d)$ be a uniformly locally finite metric space. For each $n\geq 0$, define the abelian group of uniformly finite $n$-chains $C_{n}^{\mathrm{uf}}(X;\mathbb{Z})$ to consist of all bounded, controlled-propagation functions $f\colon X^{n+1}\to\mathbb{Z}$. For $n\geq 0$ and $f\in C_{n+1}^{\mathrm{uf}}(X;\mathbb{Z})$, define the boundary operator
$\partial_{n+1}(f)(x_{0},\ldots,x_{n})=\sum_{i=0}^{n+1}(-1)^{i}\sum_{y\in X}f(x_{0},\ldots,x_{i-1},y,x_{i},\ldots,x_{n}),$
where, in the inner sum, the variable $y$ is inserted in the $i$-th coordinate. The map $\partial_{n+1}$ is a well-defined group homomorphism
$\partial_{n+1}\colon C_{n+1}^{\mathrm{uf}}(X;\mathbb{Z})\longrightarrow C_{n}^{\mathrm{uf}}(X;\mathbb{Z}),$
and the identities $\partial_{n}\circ\partial_{n+1}=0$ hold for all $n\geq 1$. Consequently, the collection $\bigl(C_{*}^{\mathrm{uf}}(X;\mathbb{Z}),\partial_{*}\bigr)$ forms a chain complex of abelian groups, which we call the uniformly finite chain complex of $(X,d)$.
Definition 1.10.
Let $(X,d)$ be a uniformly locally finite metric space. The $n$th uniformly finite homology group of $(X,d)$ with integer [^8] One can similarly define the uniformly finite chain complex and uniformly finite homology with coefficients in $\mathbb{R}$. coefficients, denoted $H_{n}^{\mathrm{uf}}(X;\mathbb{Z})$, is defined to be the $n$th homology group of the uniformly finite chain complex of $(X,d)$.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $f\colon X\to Y$ be a uniformly finite-to-one coarse map. Define a homomorphism
$f_{*}\colon C_{n}^{\mathrm{uf}}(X;\mathbb{Z})\longrightarrow C_{n}^{\mathrm{uf}}(Y;\mathbb{Z}),\quad f_{*}(g)(\bar{y})=\sum_{f(\bar{x})=\bar{y}}g(\bar{x}),\,\text{ for }\,\bar{y}\in Y^{n+1}.$
It is straightforward to check that $f_{*}$ is a well-defined group homomorphism and that it commutes with the boundary operators. Consequently, it induces a homomorphism in homology,
$H_{n}^{\mathrm{uf}}(f)\colon H_{n}^{\mathrm{uf}}(X;\mathbb{Z})\longrightarrow H_{n}^{\mathrm{uf}}(Y;\mathbb{Z}).$
Moreover, if two uniformly finite-to-one coarse maps $f,g\colon(X,d)\to(Y,\partial)$ are close, then the induced maps in uniformly finite homology coincide. In particular, for each $n\geq 0$, uniformly finite homology defines a functor
$H_{n}^{\mathrm{uf}}\colon\mathbf{Coarse}\longrightarrow\mathbf{Ab}$
from the coarse category to the category of abelian groups.
In this paper, we shall be primarily concerned with the case $n=0$. Recall from [20, Definition 10.21] that a partial translation on a metric space $(X,d)$ is a partially defined bijection
$t\colon X\dashrightarrow X\qquad\text{that satisfies}\quad\sup_{x\in\operatorname{dom}(t)}d(x,t(x))<\infty.$
Note that the above condition holds if and only if the indicator function of $\mathbbm{1}_{\operatorname{Graph}(t)}$ belongs to $C_{1}^{\text{uf}}(X;\mathbb{Z})$. From the coarse-geometric perspective, a partial translation exhibits its domain and range as identical subspaces of $(X,d)$. Given a partial translation $t\colon X\dashrightarrow X$, the indicator function of its graph,
$\mathbbm{1}_{\operatorname{Graph}(t)}\colon X\times X\longrightarrow\mathbb{Z},$
is a controlled propagation function in the sense of Definition 1.9. A direct computation shows that
$\partial_{1}\!\left(\mathbbm{1}_{\operatorname{Graph}(t)}\right)=\mathbbm{1}_{\operatorname{dom}(t)}-\mathbbm{1}_{\operatorname{ran}(t)}.$
In particular, if two subsets of $X$ are related by a partial translation, then they define the same class in the uniformly finite homology group $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$.
Remark 1.11.
Let $(X,d)$ be a uniformly locally finite metric space. The group $C_{1}^{\mathrm{uf}}(X;\mathbb{Z})$ is generated by indicator functions of entourages $E\subset X\times X$. The argument of [20, Lemma 4.10] shows that for any such entourage $E$ there exist partial translations $t_{1},\ldots,t_{n}\colon X\dashrightarrow X$ whose graphs are pairwise disjoint subsets of $E$ and satisfy
$E=\bigsqcup_{i=1}^{n}\operatorname{Graph}(t_{i}).$
In particular, the indicator function of $E$ decomposes as a finite sum of indicator functions of the graphs $\operatorname{Graph}(t_{i})$. Consequently, the group $C_{1}^{\mathrm{uf}}(X;\mathbb{Z})$ is generated by indicator functions of graphs of partial translations of $(X,d)$.
The homology class of $\mathbbm{1}_{X}$ in $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is referred to as the fundamental class of $(X,d)$. It plays a central role in determining which coarse equivalences are close to bijective ones. This is made precise by the following theorem of Whyte.
Theorem 1.12 ([29, Theorem A]).
Let $(X,d)$, $(Y,\partial)$ be uniformly locally finite metric spaces, and let $f\colon X\to Y$ be a coarse equivalence. The following are equivalent:
  1. (1)
    There exists a bijective coarse equivalence $\tilde{f}\colon X\to Y$ that is close to $f$;
  2. (2)
    $H^{\mathrm{uf}}_{0}(f)\big([\mathbbm{1}_{X}]\big)=[\mathbbm{1}_{Y}]$.
It is straightforward to verify that for a finite metric space $(X,d)$, one has $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\cong\mathbb{Z}$. In contrast, the 0th uniformly finite homology group is often highly non-trivial for infinite spaces. For example, $H_{0}^{\mathrm{uf}}(\mathbb{Z};\mathbb{Z})$ is an infinite-dimensional real vector space.
Definition 1.13.
A uniformly locally finite metric space $(X,d)$ is said to be coarsely connected if there exists $R\geq 0$ such that for any $x,y\in X$ there are points $x_{0},x_{1},\ldots,x_{n+1}\in X$ satisfying
$x=x_{0},\qquad y=x_{n+1},\qquad d(x_{i},x_{i+1})\leq R,\,\text{ for all }0\leq i\leq n.$
The following theorem shows that, for infinite coarsely connected metric spaces, the 0th uniformly finite homology group is either trivial or a real vector space.
Theorem 1.14 ([29, Corollary 3.1 and Corollary 3.2]).
Let $(X,d)$ be an infinite uniformly locally finite metric space. Then the extension of scalars
$i\colon H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\longrightarrow H_{0}^{\mathrm{uf}}(X;\mathbb{R})$
is an injective group homomorphism. Moreover, if $(X,d)$ is coarsely connected, then $i$ is an isomorphism.
In particular, if $(X,d)$ is an amenable, coarsely connected, uniformly locally finite metric space, then $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is a non-zero real vector space; see [29, Theorem D]. It also follows from the above result that the 0th uniformly finite homology group is always torsion-free.
Uniformly finite homology provides information about the structure of the $K_{0}$-group of the uniform Roe algebra. The $K_{0}$-group of $\ell^{\infty}(X)$ identifies with $\ell^{\infty}(X,\mathbb{Z})$, which, by definition, coincides with the group of uniformly finite 0-chains $C_{0}^{\mathrm{uf}}(X;\mathbb{Z})$. The canonical inclusion $i\colon\ell^{\infty}(X)\hookrightarrow C^{*}_{u}(X)$ induces a group homomorphism
$K_{0}(i)\colon\ell^{\infty}(X,\mathbb{Z})\longrightarrow K_{0}(C^{*}_{u}(X)).$
It is straightforward to verify that the composition $K_{0}(i)\circ\partial_{1}$ vanishes: By Remark 1.11, the indicator functions of partial translations generate $C_{1}^{\mathrm{uf}}(X;\mathbb{Z})$; therefore, it suffices to show that for any partial translation $t\colon X\dashrightarrow X$ one has
(1) $K_{0}(i)(\mathbbm{1}_{\operatorname{dom}(t)})-K_{0}(i)(\mathbbm{1}_{\operatorname{ran}(t)})=K_{0}(i)\circ\partial_{1}(\mathbbm{1}_{\operatorname{Graph}(t)})=0.$
It is enough to find a partial isometry in the uniform Roe algebra that intertwines the two projections. Let $v\in\mathcal{B}(\ell^{2}(X))$ be the partial isometry defined by
$v(\delta_{x})=\begin{cases}\delta_{t(x)}&\text{if }x\in\operatorname{dom}(t),\\ 0&\text{otherwise.}\end{cases}$
Then $vv^{*}=\mathbbm{1}_{\operatorname{ran}(t)}$ and $v^{*}v=\mathbbm{1}_{\operatorname{dom}(t)}$. Since $t$ is a partial translation, the partial isometry $v$ has finite propagation and therefore belongs to the uniform Roe algebra of $(X,d)$. Consequently, (1) holds.
Definition 1.15.
Let $(X,d)$ be a uniformly locally finite metric space. The canonical map $K_{0}(i)\colon\ell^{\infty}(X,\mathbb{Z})\to K_{0}(C^{*}_{u}(X))$ descends to a group homomorphism
$\alpha_{0}\colon H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\longrightarrow K_{0}(C^{*}_{u}(X)),$
which will be referred to as the 0th comparison map.
The map $\alpha_{0}$ coincides with the 0th comparison map appearing in Conjecture Conjecture for coarse groupoids. It remains an open problem whether this map is split injective in general.
2. Uniform covering isometries
Let $H$ be a separable Hilbert space. It is well known (for instance, see [30, Section 4.3]) that, for a proper coarse map between two uniformly locally finite metric spaces $f\colon(X,d)\to(Y,\partial)$, there exists an isometry $S\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$, called a covering isometry of $f$, such that the support of $S$ is asymptotic to the graph of $f$ and $\operatorname{Ad}_{S}$ restricts to a *-homomorphism between the associated Roe algebras. Moreover, the induced map in $K$-theory depends only on $f$. However, covering isometries do not, in general, induce *-homomorphisms between uniform Roe algebras or their stabilisations. In this section, we introduce a modification of the above notion adapted to the setting of stabilised uniform Roe algebras.
Definition 2.1.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $H$ be a separable Hilbert space. Given a coarse map $f\colon(X,d)\to(Y,\partial)$, an isometry $S\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$ is called a uniform cover if the following conditions are satisfied:
  1. (1)
    There exists $R\geq 0$ such that, for any $x\in X$ and $y\in Y$ with $\|\mathbbm{1}_{y}S\mathbbm{1}_{x}\|\neq 0$, one has $d(y,f(x))\leq R$; in other words, $S$ covers $f$.
  2. (2)
    For any finite-dimensional subspace $V\subset H$, there exists a finite-dimensional subspace $W\subset H$ such that $S$ restricts to a map $\ell^{2}(X,V)\to\ell^{2}(Y,W)$.
  3. (3)
    For any finite-dimensional subspace $W\subset H$, there exists a finite-dimensional subspace $V\subset H$ such that $S^{*}$ restricts to a map $\ell^{2}(Y,W)\to\ell^{2}(X,V)$.
The first condition in the above definition asserts that $S$ is a covering isometry of $f$. The motivation for the last two conditions is that we require the isometries to induce maps between the filtrations of $C^{*}_{u}(X)\otimes\mathbb{K}$ introduced in Proposition 1.6. More precisely, for every $R\geq 0$ and a finite-dimensional subspace $V\subset H$, there should exist $S\geq 0$ and a finite dimensional subspace $W\subset H$ such that $\operatorname{Ad}_{S}$ restricts to a map
$\operatorname{Ad}_{S}\colon\mathbb{C}[X,R,V]\to\mathbb{C}[Y,S,W].$
This property is demonstrated in the proof of Lemma 2.4 below. The following lemma shows that an approximable operator – hence, in particular, of controlled propagation – which satisfies conditions $(2)$ and $(3)$ of Definition 2.1 defines an element of the multiplier algebra of $C^{*}_{u}(X)\otimes\mathbb{K}$.
Lemma 2.2.
Let $(X,d)$ be a uniformly locally finite metric space. Any approximable operator $S\in\mathcal{B}(\ell^{2}(X,H))$ that satisfies conditions (2) and (3) of Definition 2.1 belongs to the multiplier algebra of $C^{*}_{u}(X)\otimes\mathbb{K}$.
Proof.
Since $C^{*}_{u}(X)\otimes\mathbb{K}$ is a non-degenerate $C^{*}$-subalgebra of $\mathcal{B}(\ell^{2}(X,H))$, it follows from [3, II 7.3.5] that its multiplier algebra can be identified with the $C^{*}$-subalgebra of $\mathcal{B}(\ell^{2}(X,H))$ given by
$\begin{split}\mathcal{A}:=\bigl\{\,\,T\in\mathcal{B}(\ell^{2}(X,H))\,\big|&\,T(C^{*}_{u}(X)\otimes\mathbb{K})\subset C^{*}_{u}(X)\otimes\mathbb{K}\\ \text{ and }&\,(C^{*}_{u}(X)\otimes\mathbb{K})T\subset C^{*}_{u}(X)\otimes\mathbb{K}\,\,\bigr\}.\end{split}$
Let $S\in\mathcal{B}(\ell^{2}(X,H))$ be an approximable operator that satisfies conditions (2) and (3) of Definition 2.1. The $C^{*}$-algebra $C^{*}_{u}(X)\otimes\mathbb{K}$ is generated by operators of the form $b\otimes F$, where $b\in C^{*}_{u}(X)$ and $F$ is a finite-rank operator. Let $W\subset H$ be a subspace such that $S$ restricts to a map
$\ell^{2}(X,\operatorname{im}(F))\longrightarrow\ell^{2}(X,W).$
Then the operator $S(b\otimes F)$ restricts to a map
$S(b\otimes F)\colon\ell^{2}(X,H)\longrightarrow\ell^{2}(X,W),$
and is approximable. Hence, $S(b\otimes F)$ belongs to $C^{*}_{u}(X)\otimes\mathbb{K}$. Since conditions (2) and (3) of Definition 2.1 are stable under taking adjoints, the operator $(b\otimes F)S$ also belongs to $C^{*}_{u}(X)\otimes\mathbb{K}$. It follows that $S\in\mathcal{A}$. ∎
It is clear from Definition 2.1 that if $S$ is a uniform cover for a coarse map $f\colon(X,d)\to(Y,\partial)$, and $g\colon(X,d)\to(Y,\partial)$ is a coarse map close to $f$, then $S$ is also a uniform cover for $g$.
For the case of coarse equivalences, the existence of covering unitaries was established in [7]. The following lemma is a refinement of [7, Theorem 4], it establishes the existence of a uniform cover for any uniformly finite-to-one coarse map.
Lemma 2.3.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, let $f\colon(X,d)\to(Y,\partial)$ be a uniformly finite-to-one coarse map, and let $H$ be a separable Hilbert space. Then there exists a uniform cover
$S\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$
of $f$. Moreover, if $f$ is a coarse equivalence, $S$ can be chosen to be unitary.
Proof.
Without loss of generality, let $H=\ell^{2}(\mathbb{N})$. Since $f$ is uniformly finite-to-one, the quantity
$N:=\sup_{y\in Y}\lvert f^{-1}(\{y\})\rvert$
is finite. For $0\leq i\leq N$, let $Y_{i}\subset Y$ denote the set of those $y\in Y$ whose preimage under $f$ consists of precisely $i$ distinct points. In this way, we obtain a partition
$Y=\bigsqcup_{i=0}^{N}Y_{i}.$
For each $1\leq i\leq N$, fix a partition $\{A_{k}^{i}\}_{k=1}^{i}$ of $\mathbb{N}$ into $i$ pairwise disjoint subsets, each of which is in bijection with $\mathbb{N}$. Let $f_{k}^{i}\colon\mathbb{N}\to A_{k}^{i}$ be a bijection. For every $y\in\operatorname{im}(f)$, enumerate the finite set $f^{-1}(\{y\})$. Define an isometry $S$ on the canonical basis of $\ell^{2}(X)\otimes\ell^{2}(\mathbb{N})$ by setting
$S(\delta_{x_{k}}\otimes\delta_{n})=\delta_{y}\otimes\delta_{f_{k}^{i}(n)},$
where $y\in Y_{i}$ for some $i\neq 0$ and $f(x_{k})=y$. It is straightforward to verify that $\operatorname{supp}(S)=f$. Given $\zeta\in\ell^{2}(\mathbb{N})$, $y\in Y_{i}$, and $x_{k}\in f^{-1}(\{y\})$, one has
$S(\delta_{x_{k}}\otimes\zeta)=\delta_{y}\otimes\zeta(k,i),\qquad\text{where}\quad\zeta(k,i):=\sum_{n\in\mathbb{N}}\langle\zeta,\delta_{n}\rangle\,\delta_{f_{k}^{i}(n)}.$
Let $H_{\zeta}$ be the subspace of $\ell^{2}(\mathbb{N})$ generated by the vectors $\zeta(k,i)$ for $1\leq k\leq i\leq N$. This subspace is finite-dimensional, and
$S\bigl(\ell^{2}(X)\otimes\zeta\bigr)\subset\ell^{2}(Y)\otimes H_{\zeta}.$
Now let $H_{0}$ be any finite-dimensional subspace of $\ell^{2}(\mathbb{N})$, and let $H_{1}$ denote the subspace generated by $H_{\zeta}$ for all $\zeta\in H_{0}$. Then $H_{1}$ is finite-dimensional, and
$S\bigl(\ell^{2}(X)\otimes H_{0}\bigr)\subset\ell^{2}(Y)\otimes H_{1}.$
Analogously, for $\eta\in\ell^{2}(\mathbb{N})$ and $y_{0}\in Y_{i}$ with $i\neq 0$, one has
$S^{*}(\delta_{y_{0}}\otimes\eta)=\sum_{k=1}^{i}\delta_{x_{k}}\otimes\eta(k,i),\qquad\text{where }\eta(k,i):=\sum_{n\in N_{k}^{i}}\langle\eta,\delta_{n}\rangle\,\delta_{(f_{k}^{i})^{-1}(n)}.$
Moreover, for any $y\in Y_{0}$, one has $S^{*}(\delta_{y}\otimes\eta)=0$. Let $H_{\eta}$ be the subspace of $\ell^{2}(\mathbb{N})$ generated by the vectors $\eta(k,i)$ for $1\leq k\leq i\leq N$. This subspace is finite-dimensional and
$S^{*}(\ell^{2}(Y)\otimes\eta)\subset\ell^{2}(X)\otimes H_{\eta}.$
For any finite-dimensional subspace $H_{0}\subset\ell^{2}(\mathbb{N})$, let $H_{1}$ denote the subspace generated by $H_{\eta}$ for all $\eta\in H_{0}$. Then $H_{1}$ is finite-dimensional, and
$S^{*}(\ell^{2}(Y)\otimes H_{0})\subset\ell^{2}(X)\otimes H_{1}.$
It follows that the constructed isometry $S$ satisfies the definition of a covering isometry for $f$. Note that if $f$ is surjective, then $S$ is surjective; in particular, $S$ is a unitary.
Suppose that $f$ is a coarse equivalence. Then it can be decomposed into maps
[…]
where $f_{0}$ is a surjective coarse equivalence, and $i$ is the inclusion of a coarsely dense subset. Applying the construction of a covering isometry described above to the surjective coarse equivalence $f_{0}$, we obtain a unitary
$U_{f_{0}}\colon\ell^{2}(X)\otimes\ell^{2}(\mathbb{N})\to\ell^{2}(f(X))\otimes\ell^{2}(\mathbb{N}).$
Since $i\colon f(X)\to Y$ is the inclusion of a coarsely dense subset, there exists a left inverse $p\colon Y\to f(X)$, which is a surjective coarse equivalence. Applying the same construction to $p$ yields a covering unitary
$U_{p}\colon\ell^{2}(Y)\otimes\ell^{2}(\mathbb{N})\to\ell^{2}(f(X))\otimes\ell^{2}(\mathbb{N}).$
It is straightforward to verify that the unitary $U_{p}^{*}U_{f_{0}}$ satisfies Definition 2.1. ∎
Note that the proof of Lemma 2.3 involves a substantial number of choices; consequently, the resulting uniform cover is highly non-canonical. For instance, if $X$ is a singleton, then any isometry in $\mathcal{B}(H)$ covers $\operatorname{id}_{X}$. In the case of covering isometries for Roe algebras, this non-canonicity disappears at the level of $K$-theory. The following lemma shows that the same phenomenon holds for uniform covers.
Lemma 2.4.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, let $f\colon(X,d)\to(Y,\partial)$ be a uniformly finite-to-one coarse map, and let $H$ be a separable Hilbert space. Let $S\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$ be a uniform cover of $f$. Then $\operatorname{Ad}_{S}$ restricts to a *-homomorphism
$\operatorname{Ad}_{S}\colon C_{u}^{*}(X)\otimes\mathbb{K}\longrightarrow C_{u}^{*}(Y)\otimes\mathbb{K}.$
Moreover, the induced map in $K$-theory does not depend on the choice of a uniform cover for $f$.
Proof.
Let $f\colon(X,d)\to(Y,\partial)$ be a uniformly finite-to-one coarse map, and let $S\colon\ell^{2}(X,H)\to\ell^{2}(Y,H)$ be a uniform cover of $f$. Fix $R\geq 0$ and a finite-dimensional subspace $H_{0}\subset H$. Since $f$ is coarse, there exists $R_{0}\geq 0$ such that for any $x,y\in X$ satisfying $d(x,y)\leq R$, one has $d(f(x),f(y))\leq R_{0}$. By the definition of a uniform cover, there exists $R_{1}\geq 0$ such that
$\|\mathbbm{1}_{y}S\mathbbm{1}_{x}\|\neq 0\;\Longrightarrow\;d(y,f(x))\leq R_{1}.$
In particular, for any operator $T\in\mathcal{B}(\ell^{2}(X,H))$ of propagation at most $R$, the operator $STS^{*}$ has propagation at most $R_{0}+2R_{1}$. Since $S$ satisfies condition $(2)$ of Definition 2.1, there exists a finite-dimensional subspace $H_{1}\subset H$ such that $S$ restricts to an isometry from $\ell^{2}(X,H_{0})$ to $\ell^{2}(Y,H_{1})$. It follows that for any $T\in\mathcal{B}(\ell^{2}(X,H_{0}))$ the operator $STS^{*}$ belongs to $\mathcal{B}(\ell^{2}(Y,H_{1}))$. In particular, $\operatorname{Ad}_{S}$ restricts to a map
$\operatorname{Ad}_{S}\colon\mathbb{C}[X,R,H_{0}]\longrightarrow\mathbb{C}[Y,R_{0}+2R_{1},H_{1}].$
By passing to closures, it follows that $\operatorname{Ad}_{S}$ restricts to a *-homomorphism between stabilisations of uniform Roe algebras.
For the $K$-theory part, let $S_{1}$ and $S_{2}$ be two uniform covers of $f$. For convenience, set $A=M_{2}(\mathbb{C})\otimes C_{u}^{*}(Y)\otimes\mathbb{K}$. For $i=1,2$, consider the maps
$\alpha_{i}\colon C_{u}^{*}(X)\otimes\mathbb{K}\longrightarrow A,\qquad\alpha_{i}(T)=e_{ii}\otimes\operatorname{Ad}_{S_{i}}(T).$
It suffices to show that $\alpha_{1}$ and $\alpha_{2}$ agree in $K$-theory. The projections $S_{1}S_{1}^{*}$ and $S_{2}S_{2}^{*}$, as well as the partial isometries $S_{1}S_{2}^{*}$ and $S_{2}S_{1}^{*}$, satisfy conditions of Lemma 2.2; therefore, they belong to the multiplier algebra of $C_{u}^{*}(Y)\otimes\mathbb{K}$. Consider the unitary
$U=\begin{pmatrix}1-S_{1}S_{1}^{*}&S_{1}S_{2}^{*}\\ S_{2}S_{1}^{*}&1-S_{2}S_{2}^{*}\end{pmatrix}$
in the multiplier algebra of $A$. It follows that $\operatorname{Ad}_{U}$ induces the identity map on $K$-theory and intertwines $\alpha_{1}$ and $\alpha_{2}$. Therefore, $K_{*}(\alpha_{1})=K_{*}(\alpha_{2})$. ∎
In light of Lemma 2.4, for a uniformly finite-to-one coarse map $f$, we denote by $K_{*}(f)$ the map $K_{*}(\operatorname{Ad}_{S})$ for some (any) uniform cover $S$. Since uniform covers depend only on the closeness class of a coarse map, the homomorphism $K_{*}(f)$ depends only on the closeness class of $f$. It follows that, for $i=1,2$, there are well-defined functors
$K_{i}\colon\mathbf{Coarse}\longrightarrow\mathbf{Ab},\qquad(X,d)\longmapsto K_{i}\bigl(C_{u}^{*}(X)\bigr),\qquad f\longmapsto K_{i}(f),$
from the coarse category to the category of abelian groups. Another family of functors from the coarse category to abelian groups is given by uniformly finite homology (see Definition 1.10). Recall from Definition 1.15 that there is a map
$\alpha_{0}\colon H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\longrightarrow K_{0}\bigl(C_{u}^{*}(X)\bigr),$
called the 0th comparison map. The following lemma asserts that $\alpha_{0}$ is a natural transformation from 0th uniformly finite homology to $K_{0}$.
Lemma 2.5.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and denote by $\alpha_{0}^{X}$ and $\alpha_{0}^{Y}$ the associated 0th comparison maps. Let $f\colon(X,d)\to(Y,\partial)$ be a uniformly finite-to-one coarse map. Then the following diagram commutes:
$H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$$K_{0}(C_{u}^{*}(X))$$H_{0}^{\mathrm{uf}}(Y;\mathbb{Z})$$K_{0}(C_{u}^{*}(Y)).$$\scriptstyle{\alpha_{0}^{X}}$$\scriptstyle{H_{0}^{\mathrm{uf}}(f)}$$\scriptstyle{K_{0}(f)}$$\scriptstyle{\alpha_{0}^{Y}}$
Proof.
Suppose that $f\colon(X,d)\to(Y,\partial)$ is an injective coarse map. Define an isometry
$S\colon\ell^{2}(X)\to\ell^{2}(Y),\qquad S(\delta_{x}):=\delta_{f(x)}.$
By definition, $\widetilde{S}:=S\otimes\operatorname{id}_{H}$ is a uniform cover of $f$; therefore, the maps $K_{0}(f)$ and $K_{0}(\operatorname{Ad}_{S})$ coincide. Moreover, the *-homomorphism $\operatorname{Ad}_{S}$ restricts to a *-homomorphism between the canonical Cartan subalgebras $\ell^{\infty}(X)$ and $\ell^{\infty}(Y)$ of $C^{*}_{u}(X)$ and $C^{*}_{u}(Y)$, respectively. In particular, the following diagram commutes:
$K_{0}(\ell^{\infty}(X))$$K_{0}(C^{*}_{u}(X))$$K_{0}(\ell^{\infty}(Y))$$K_{0}(C^{*}_{u}(Y)),$$\scriptstyle{K_{0}(i_{X})}$$\scriptstyle{K_{0}(\operatorname{Ad}_{S})}$$\scriptstyle{K_{0}(f)}$$\scriptstyle{K_{0}(i_{Y})}$
where $i_{X}$ and $i_{Y}$ denote the canonical inclusions. For a uniformly locally finite metric space $(Z,d_{Z})$, let $\pi_{Z}$ be the canonical projection from $K_{0}(\ell^{\infty}(Z))$ onto $H_{0}^{\mathrm{uf}}(Z;\mathbb{Z})$. Since $f$ is injective, a direct computation shows that, for any subset $A\subseteq X$, one has $\operatorname{Ad}_{S}(\mathbbm{1}_{A})=\mathbbm{1}_{f(A)}$ and $f_{*}(\mathbbm{1}_{A})=\mathbbm{1}_{f(A)}$. The following diagram commutes:
$K_{0}(\ell^{\infty}(X))$$H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$$K_{0}(\ell^{\infty}(Y))$$H_{0}^{\mathrm{uf}}(Y;\mathbb{Z}).$$\scriptstyle{\pi_{X}}$$\scriptstyle{K_{0}(\operatorname{Ad}_{S})}$$\scriptstyle{H_{0}^{\mathrm{uf}}(f)}$$\scriptstyle{\pi_{Y}}$
Finally, since the map on $K_{0}$ induced by the inclusion $\ell^{\infty}(X)\hookrightarrow C^{*}_{u}(X)$ factors through the 0th comparison map, the diagram appearing in the statement commutes.
For the general case, let $f\colon(X,d)\to(Y,\partial)$ is a uniformly finite-to-one coarse map, and set
$n:=\sup_{y\in Y}\lvert f^{-1}(y)\rvert$.
Let $Y^{(n)}$ denote the $n$th doubling of $Y$, and let $j\colon Y\to Y^{(n)}$ be the inlcusion of $Y$ onto $Y\times\{1\}\subset Y^{(n)}$. The inclusion $j$ is a coarse equivalence. For every $y\in Y$ whose preimage has $i\neq 0$ points, fix an enumeration $x_{1}^{y},\ldots,x_{i}^{y}$ of $f^{-1}(\{y\})$. Define a map
$\tilde{f}\colon X\to Y^{(n)},\qquad x_{i}^{f(x)}\mapsto(f(x),i).$
It is straightforward to verify that $\tilde{f}$ is injective, and it is close to $j\circ f$. In particular, one has the equalities
$H_{0}^{\mathrm{uf}}(\tilde{f})=H_{0}^{\mathrm{uf}}(j_{1})\circ H_{0}^{\mathrm{uf}}(f),\qquad K_{0}(\tilde{f})=K_{0}(j_{1})\circ K_{0}(f).$
Since $\tilde{f}$ and $j$ are injective, the diagrams appearing in the statement commute for $\tilde{f}$ and $j$ by the first part of the proof. Since $j$ is a coarse equivalence, the maps $H_{0}^{\mathrm{uf}}(j)$ and $K_{0}(j)$ are isomorphisms. The following diagram commutes:
$H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$$K_{0}(C^{*}_{u}(X))$$H_{0}^{\mathrm{uf}}(Y^{(n)};\mathbb{Z})$$K_{0}(C^{*}_{u}(Y^{(n)}))$$H_{0}^{\mathrm{uf}}(Y;\mathbb{Z})$$K_{0}(C^{*}_{u}(Y)).$$\scriptstyle{H_{0}^{\mathrm{uf}}(f)}$$\scriptstyle{\alpha_{0}^{X}}$$\scriptstyle{H_{0}^{\mathrm{uf}}(\tilde{f})}$$\scriptstyle{K_{0}(\tilde{f})}$$\scriptstyle{K_{0}(f)}$$\scriptstyle{H_{0}^{\mathrm{uf}}(j_{1})^{-1}}$$\scriptstyle{\alpha_{0}^{Y^{(n)}}}$$\scriptstyle{K_{0}(j_{1})^{-1}}$$\scriptstyle{\alpha_{0}^{Y}}$
In particular, the diagram appearing in the statement commutes. ∎
There is a more conceptual explanation of Lemma 2.5, which we briefly outline without entering into technical details. Recall from Section 1.1 that to each uniformly locally finite metric space $(X,d)$ one can associate a principal ample groupoid $\mathcal{G}(X,d)$, called the coarse groupoid of $(X,d)$. Following Proposition 2.3 of [23], any uniformly finite-to-one coarse map
$f\colon(X,d)\longrightarrow(Y,\partial)$
induces a uniformly locally finite coarse structure $\mathcal{E}_{f}$ on the disjoint union $X\sqcup Y$, whose restriction to $Y$ coincides with the original coarse structure of $Y$, and whose restriction to $X$ contains the original coarse structure of $X$. Consequently, at the level of coarse groupoids, one obtains étale groupoids homomorphisms
$\mathcal{G}(X,d)$$\mathcal{G}(X\sqcup Y,\mathcal{E}_{f})$$\mathcal{G}(Y,\partial),$$\scriptstyle{i_{X}}$$\scriptstyle{i_{Y}}$
where the right-hand arrow is a Morita equivalence of groupoids. Passing to groupoid homology or to the $K$-theory of reduced groupoid $C^{*}$-algebras yields group homomorphisms
$F(f)\colon F(\mathcal{G}(X,d))\longrightarrow F(\mathcal{G}(Y,\partial)),\qquad F(f)=F(i_{Y})^{-1}\circ F(i_{X}),$
where $F$ denotes either groupoid homology or the $K$-theory of reduced groupoid $C^{*}$-algebras. The 0th comparison map for ample groupoids is a natural transformation; therefore, the following diagram commutes:
$H_{0}(\mathcal{G}(X,d);\mathbb{Z})$$K_{0}(C_{r}^{*}(\mathcal{G}(X,d)))$$H_{0}(\mathcal{G}(Y,\partial);\mathbb{Z})$$K_{0}(C_{r}^{*}(\mathcal{G}(Y,\partial))).$$\scriptstyle{\alpha_{0}}$$\scriptstyle{H_{0}(f)}$$\scriptstyle{K_{0}(f)}$$\scriptstyle{\alpha_{0}}$
Moreover, by [5, Theorem G], there is a natural isomorphism between the homology of coarse groupoids and uniformly finite homology. As the uniform Roe algebra is naturally isomorphic to the reduced groupoid $C^{*}$-algebra of the associated coarse groupoid, the commutativity of the diagram in Lemma 2.5 follows.
3. Rigidity of uniform Roe algebras versus injectivity of $\alpha_{0}$
In this section, we establish a close connection between the injectivity of the 0th comparison map for coarse groupoids and the bijective rigidity problem. In contrast to arguments involving a single space, our approach necessarily requires the consideration of subspaces and doublings.
Definition 3.1.
A collection $\mathcal{C}$ of uniformly locally finite metric spaces is called admissible if it is closed under taking arbitrary doublings and subspaces.
For example, let $\mathcal{C}(\mathrm{A})$ denote the collection of all uniformly locally finite metric spaces that satisfy Property A, and let $\mathcal{C}(\mathrm{CE})$ denote the collection of all uniformly locally finite metric spaces that admit a coarse embedding into a Hilbert space. Since both properties are preserved under taking subspaces and under passage to coarsely equivalent spaces, the collections $\mathcal{C}(\mathrm{A})$ and $\mathcal{C}(\mathrm{CE})$ are admissible. Clearly, the collection of all uniformly locally finite metric spaces is also admissible. The following lemma is the main result of this section.
Lemma 3.2 (Lemma A).
Let $\mathcal{C}$ be an admissible collection of uniformly locally finite metric spaces. The following statements are equivalent:
  1. (1)
    For every uniformly locally finite metric space $(X,d)$ in $\mathcal{C}$, the 0th comparison map $\alpha_{0}\colon H^{\mathrm{uf}}_{0}(X;\mathbb{Z})\to K_{0}\bigl(C^{*}_{u}(X)\bigr)$ is injective;
  2. (2)
    For any uniformly locally finite metric spaces $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ such that $C^{*}_{u}(Y)\cong C^{*}_{u}(X)$, the coarse equivalence given by Theorem 1.7 is close to a bijective coarse equivalence.
Note that, for ample groupoids, the injectivity of $\alpha_{0}$ is invariant under Morita equivalence; hence, the injectivity of $\alpha_{0}$ is preserved by passing to coarsely equivalent spaces. In particular, the 0th comparison map for $(X,d)$ is injective if and only if it is injective for any (some) doubling of $(X,d)$. On the other hand, it is not evident whether injectivity passes to subspaces. Before proceeding with the proof of Lemma 3.2, we need to establish a connection between the $K$-theory of an isomorphism $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$, and the $K$-theory of the coarse equivalence $f_{\Phi}$ induced by Theorem 1.7.
Proposition 3.3.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$ be an isomorphism. Denote by $f_{\Phi}\colon(X,d)\to(Y,\partial)$ the coarse equivalence provided by Theorem 1.7. Then the induced maps in $K$-theory, $K_{*}(f_{\Phi})$ and $K_{*}(\Phi)$, coincide.
Proof.
Let $U$ be a uniform cover for $f_{\Phi}$. By [24, Lemma 3.1], any isomorphism between uniform Roe algebras is spatially implemented; hence, there exists a unitary operator $V\colon\ell^{2}(X)\to\ell^{2}(Y)$ such that $\Phi=\operatorname{Ad}_{V}$. Consider the unitary operator
$\widetilde{U}:=U^{*}(V\otimes\operatorname{id}_{H}).$
Note that $\widetilde{U}$ satisfies conditions (2) and (3) of Definition 2.1. By Theorem 1.8, the unitary $V\otimes\operatorname{id}_{H}$ is a norm limit of a net $\{T_{\lambda}\}_{\lambda}$ of operators supported on $f_{\Phi}$. It follows that $\widetilde{U}$ is a norm limit of operators supported on $\operatorname{id}_{X}$; in particular, $\widetilde{U}$ is approximable. By Lemma 2.2, the unitary $\widetilde{U}$ belongs to the multiplier algebra of $C^{*}_{u}(X)\otimes\mathbb{K}$ and therefore induces the identity map in $K$-theory. Consequently,
$\operatorname{id}=K_{*}(\operatorname{Ad}_{\widetilde{U}})=K_{*}(\Phi)^{-1}\circ K_{*}(f_{\Phi}),$
and hence the maps $K_{*}(\Phi)$ and $K_{*}(f_{\Phi})$ coincide. ∎
Remark 3.4.
In the proof of Proposition 3.3, we showed that there exists a unitary $\tilde{U}$ in the multiplier algebra of $C^{*}_{u}(X)\otimes\mathbb{K}$ which intertwines the *-isomorphisms $\Phi\otimes\operatorname{id}_{H}$ and $\operatorname{Ad}_{U}$, where $U$ is a uniform cover of $f_{\Phi}$. Recall from [18, Theorem 2.5] that for any unital $C^{*}$-algebra $A$, the unitary group of the multiplier algebra of $A\otimes\mathbb{K}$ is contractible. Consequently, there exists a continuous path of unitaries $u_{t}\in\mathcal{M}(C^{*}_{u}(X)\otimes\mathbb{K})$ such that $u_{0}=\tilde{U}$ and $u_{1}=1$. In particular, the *-isomorphisms $\Phi\otimes\operatorname{id}_{H}$ and $\operatorname{Ad}_{U}$ are homotopic.
By combining Proposition 3.3 and Lemma 2.5, we deduce that for any *-isomorphism of uniform Roe algebras $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$ the following diagram commutes:
(2) $H_{0}^{\text{uf}}(X;\mathbb{Z})$$K_{0}(C^{*}_{u}(X))$$H_{0}^{\text{uf}}(Y;\mathbb{Z})$$K_{0}(C^{*}_{u}(Y)),$$\scriptstyle{\alpha_{0}^{X}}$$\scriptstyle{H_{0}^{\text{uf}}(f_{\Phi})}$$\scriptstyle{K_{0}(\Phi)}$$\scriptstyle{\alpha_{0}^{Y}}$
where $f_{\Phi}$ denotes the coarse equivalence induced by $\Phi$ via Theorem 1.7.
Proof of Lemma 3.2.
Let $\mathcal{C}$ be an admissible collection of uniformly locally finite metric spaces such that, for every $(X,d)\in\mathcal{C}$, the comparison map
$\alpha_{0}^{X}\colon H_{0}^{\text{uf}}(X;\mathbb{Z})\to K_{0}(C^{*}_{u}(X))$
is injective. Suppose that $(X,d)$ and $(Y,\partial)$ belong to $\mathcal{C}$ and $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$ is a *-isomorphism. By the commutativity of (2), we obtain
$\alpha_{0}^{Y}\circ H_{0}^{\text{uf}}(f_{\Phi})([\mathbbm{1}_{X}])=K_{0}(\Phi)\circ\alpha_{0}^{X}([\mathbbm{1}_{X}]),$
and the right-hand side is exactly the $K_{0}$-class of $\mathbbm{1}_{Y}$. Since $\alpha_{0}^{Y}$ is injective, it follows that $H_{0}^{\text{uf}}(f_{\Phi})$ preserves the fundamental class. By Theorem 1.12, there exists a bijective coarse equivalence that is close to $f_{\Phi}$. Consequently, the coarse equivalence given by Theorem 1.7 is close to a bijective coarse equivalence.
Conversely, let $\mathcal{C}$ be an admissible collection of uniformly locally finite metric spaces such that for any $(X,d)$ and $(Y,\partial)$ in $\mathcal{C}$ satisfing $C^{*}_{u}(Y)\cong C^{*}_{u}(X)$, the coarse equivalence given by Theorem 1.7 is close to a bijective coarse equivalence. Let $(X,d)\in\mathcal{C}$, and let $\beta\in\ker(\alpha_{0}^{X})$. By passing, if necessary, to a doubling of $(X,d)$, for some $n\geq 1$ the class $\beta$ is given by
$\beta=[\mathbbm{1}_{B}]-[\mathbbm{1}_{A}],$
for some subsets $A,B\subset X^{(n)}$. Since $\beta$ lies in the kernel of $\alpha_{0}^{X}$, the projections $\mathbbm{1}_{A}$ and $\mathbbm{1}_{B}$ define the same class in $K_{0}(C^{*}_{u}(X^{(n)}))$. Consequently, there exist $\ell,k\in\mathbb{N}$ and a partial isometry $s$ in $C^{*}_{u}(X^{(n)})\otimes M_{\ell+k+1}(\mathbb{C})$ such that
$\mathbbm{1}_{A}\oplus\mathbbm{1}_{X^{(n)}}^{\oplus\ell}\oplus 0^{\oplus k}=ss^{*}\sim s^{*}s=\mathbbm{1}_{B}\oplus\mathbbm{1}_{X^{(n)}}^{\oplus\ell}\oplus 0^{\oplus k}.$
It follows that $\Phi:=\operatorname{Ad}_{s}$ induces an isomorphism between the uniform Roe algebras of $A\sqcup X^{(n+\ell)}$ and $B\sqcup X^{(n+\ell)}$. By Theorem 1.7, the isomorphism $\Phi$ gives rise to a coarse equivalence
$f_{\Phi}\colon A\sqcup X^{(n+\ell)}\longrightarrow B\sqcup X^{(n+\ell)}.$
Recall that, up to closeness, for some (equivalently, any) $\delta\in(0,1)$, there exist constants $R,P\geq 0$ such that the relation $f_{\Phi}$ has the following form:
$f_{\Phi}=\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|>\delta\}.$
Since $s$ is approximable, for every $\varepsilon>0$ there exists a controlled propagation operator $t$ such that $\|s-t\|<\varepsilon$. Choose $\varepsilon>0$ with $\varepsilon<\delta$. If $C,D\subset X$ satisfy $\|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|>\delta$, then
$\delta<\|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|\leq\|\mathbbm{1}_{C}(s-t)\mathbbm{1}_{D}\|+\|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|<\varepsilon+\|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|.$
In particular, $\|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|>\delta-\varepsilon$. Consequently, we obtain a chain of inclusions
$\begin{split}f_{\Phi}=\;&\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}s\mathbbm{1}_{D}\|>\delta\}\\ \subset\;&\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|>\delta-\varepsilon\}\\ \subset\;&\bigcup\{C\times D\mid\operatorname{diam}(C)\leq R,\ \operatorname{diam}(D)\leq P,\ \|\mathbbm{1}_{C}t\mathbbm{1}_{D}\|\neq 0\}\\ \subset\;&E_{R}\circ\operatorname{supp}(t)\circ E_{P}.\end{split}$
Since $t$ has controlled propagation, its support is an entourage. As entourages are closed under composition, the latter set is also an entourage. In particular, the relation $f_{\Phi}$ is contained in an entourage; therefore, it is close to the identity map on $X$. By assumption, $f_{\Phi}$ can be chosen to be bijective; in particular, $f_{\Phi}$ is a partial translation. Hence the indicator functions of $A\sqcup X^{(n+\ell)}$ and $B\sqcup X^{(n+\ell)}$ define the same class in $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$. By substracting the homology class of the indicator function of $X^{(n+\ell)}$, we conclude that $\mathbbm{1}_{A}]=[\mathbbm{1}_{B}]$, and therefore $\beta=0$. This shows that $\alpha_{0}^{X}$ is injective. ∎
Note that the proof of the first implication does not rely on the assumption that $\mathcal{C}$ is an admissible collection. Indeed, the argument does not involve passing to subspaces or to doublings of the underlying metric spaces. Consequently, the conclusion of the first implication holds for a single space.
Corollary 3.5.
Let $(X,d)$ be a uniformly locally finite metric space for which the 0th comparison map $\alpha_{0}^{X}$ is injective. Then, for any uniformly locally finite metric space $(Y,\partial)$ whose uniform Roe algebra is isomorphic to $C^{*}_{u}(X)$, the coarse equivalence provided by Theorem 1.7 is close to a bijective coarse equivalence.
Proof.
Let $(X,d)$ and $(Y,\partial)$ be uniformly locally finite metric spaces, and let $\Phi\colon C^{*}_{u}(X)\to C^{*}_{u}(Y)$ be an isomorphism. Denote by $f_{\Phi}$ the coarse equivalence provided by Theorem 1.7. Assume that the 0th comparison map $\alpha_{0}^{X}$ is injective. By Lemma 2.5, since $H_{0}^{\mathrm{uf}}(f_{\Phi})$ and $K_{0}(f_{\Phi})$ are isomorphisms, it follows that the 0th comparison map $\alpha_{0}^{Y}$ is also injective. Consequently, by the proof of Lemma 3.2, the coarse equivalence $f_{\Phi}$ is close to a bijective coarse equivalence. ∎
Remark 3.6.
In light of Theorem 1.14, the injectivity of the 0th comparison map is equivalent to the injectivity of the rational comparison map
$\alpha_{0}^{X}\otimes\operatorname{id}_{\mathbb{Q}}\colon H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Q}\longrightarrow K_{0}(C^{*}_{u}(X))\otimes_{\mathbb{Z}}\mathbb{Q}.$
Indeed, since $\mathbb{Q}$ is a torsion-free abelian group, it is projective; consequently, the functor $\otimes_{\mathbb{Z}}\mathbb{Q}$ preserves injective group homomorphisms. In particular, if $\alpha_{0}^{X}$ is injective, then so is $\alpha_{0}^{X}\otimes\operatorname{id}_{\mathbb{Q}}$. Conversely, for any torsion-free abelian group $A$, the canonical map
$i_{A}\colon A\longrightarrow A\otimes_{\mathbb{Z}}\mathbb{Q},\qquad a\longmapsto a\otimes 1,$
is injective. By Theorem 1.14, for any uniformly locally finite metric space $(X,d)$, the group $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is torsion-free. Consider the following commutative diagram:
$H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Q}$$K_{0}(C^{*}_{u}(X))\otimes_{\mathbb{Z}}\mathbb{Q}$$H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$$K_{0}(C^{*}_{u}(X)).$$\scriptstyle{\alpha_{0}^{X}\otimes\operatorname{id}_{\mathbb{Q}}}$$\scriptstyle{\alpha_{0}^{X}}$$\scriptstyle{i}$$\scriptstyle{i}$
It follows that injectivity of $\alpha_{0}^{X}\otimes\operatorname{id}_{\mathbb{Q}}$ implies injectivity of $\alpha_{0}^{X}$. Consequently, in Lemma 3.2 and its corollaries, it suffices to assume rational injectivity of the 0th comparison map.
4. Consequences
In [28, Theorem E] it is shown that the class $\mathcal{C}(\mathrm{A})$ of uniformly locally finite metric spaces satisfying property A (equivalently, the class of amenable coarse groupoids) satisfies bijective rigidity. By applying Lemma 3.2, we obtain that the 0th comparison map is injective for all uniformly locally finite spaces that satisfy Property A. The authors of [5] (see Corollary C) proved that for any ample, principal, second-countable, locally compact, Hausdorff groupoid $\mathcal{G}$ with dynamic asymptotic dimension at most $2$, such that $H_{2}(\mathcal{G};\mathbb{Z})$ is finitely generated, there are isomorphisms
$K_{0}(C^{*}_{r}(\mathcal{G}))\cong H_{0}(\mathcal{G};\mathbb{Z})\oplus H_{2}(\mathcal{G};\mathbb{Z}),\qquad K_{1}(C^{*}_{r}(\mathcal{G}))\cong H_{1}(\mathcal{G};\mathbb{Z}).$
They also demonstrated that the result applies to coarse groupoids. For coarse groupoids, the dynamical asymptotic dimension coincides with the asymptotic dimension of the underlying metric space. Consequently, the above isomorphisms hold for all uniformly locally finite metric spaces of asymptotic dimension at most $2$ with finitely generated $H_{2}^{\text{uf}}(X;\mathbb{Z})$. Since every uniformly locally finite metric space of finite asymptotic dimension has property A, our results may be viewed as a partial generalisation of the results of [5]. Recently, independently of this work, Vignati showed in [27] that bijective rigidity holds for all uniformly locally finite metric spaces (see Theorem 0.2). Moreover, the bijective coarse equivalence constructed therein is close to the coarse equivalence provided by Theorem 1.7. We obtain the following theorem.
Theorem 4.1 (Theorem B).
Let $(X,d)$ be a uniformly locally finite metric space, and let $\mathcal{G}$ denote its coarse groupoid. Then the comparison map
$\alpha_{0}\colon H_{0}(\mathcal{G};\mathbb{Z})\to K_{0}(C^{*}_{r}(\mathcal{G}))$
is injective. Moreover, if $(X,d)$ is coarsely connected, then $\alpha_{0}$ is split injective.
Proof.
By [27], the second assertion of Lemma 3.2 holds; therefore, for any uniformly locally finite metric space $(X,d)$, the comparison map $\alpha_{0}$ is injective. Suppose now that $(X,d)$ is coarsely connected. By Theorem 1.14, the canonical map
$H_{0}^{\mathrm{uf}}(X;\mathbb{Z})\longrightarrow H_{0}^{\mathrm{uf}}(X;\mathbb{R})$
is an isomorphism. In particular, $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ carries the structure of a real vector space. It follows that any injective homomorphism from $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ splits. ∎
Let $G$ be a finitely generated group equipped with the word-length metric. By [10, Proposition A.10], its 0th uniformly finite homology group canonically identifies with the group homology of $G$ with coefficients in $\ell^{\infty}(G,\mathbb{R})$. Since Cayley graphs of finitely generated groups are coarsely connected, we obtain the following application.
Corollary 4.2 (Corollary C).
Let $G$ be a finitely generated group equipped with the word-length metric. The canonical map
$\alpha_{0}^{G}\colon H_{0}\bigl(G;\ell^{\infty}(G,\mathbb{R})\bigr)\longrightarrow K_{0}\bigl(\ell^{\infty}(G)\rtimes_{r}G\bigr)$
is split-injective.
Corollary 4.2 relies on the fact that real vector spaces are injective objects in the category of abelian groups. In general, however, $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ need not be injective. For example, if $X$ is finite, then $H_{0}^{\mathrm{uf}}(X;\mathbb{Z})$ is isomorphic to $\mathbb{Z}$. As another example, consider the subset $\mathbb{N}_{s}:=\{n^{2}\mid n\in\mathbb{N}\}$ of $\mathbb{N}$ equipped with the induced metric. This space is sometimes referred to as the sparse natural numbers. One readily verifies that
$H_{0}^{\mathrm{uf}}(\mathbb{N}_{s};\mathbb{Z})\cong\ell^{\infty}(\mathbb{N},\mathbb{Z})/\{f\in c_{0}(\mathbb{N},\mathbb{Z})\mid\sum_{n\in\mathbb{N}}f(n)=0\}.$
The latter group is not divisible and therefore is not injective. Note that, for any abelian group $A$, the group $A\otimes_{\mathbb{Z}}\mathbb{Q}$ is a $\mathbb{Q}$-vector space and therefore injective. It is an immediate consequence of Theorem 4.1 and Remark 3.6 that the rational comparison map
$\alpha_{0}\otimes\operatorname{id}_{\mathbb{Q}}\colon H_{0}(\mathcal{G};\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Q}\to K_{0}(C^{*}_{r}(\mathcal{G}))\otimes_{\mathbb{Z}}\mathbb{Q}$
is split-injective. In [19], the authors show that for any second-countable, locally compact, Hausdorff, ample groupoid $\mathcal{G}$ with torsion-free stabilisers, there is a group isomorphism
(3) $K_{*}^{\mathrm{top}}(\mathcal{G},C_{0}(\mathcal{G}^{(0)}))\otimes_{\mathbb{Z}}\mathbb{Q}\longrightarrow\bigoplus_{k\geq 0}H_{*+2k}(\mathcal{G};\mathbb{Z})\otimes_{\mathbb{Z}}\mathbb{Q},$
where the left-hand side denotes the topological $K$-theory of $\mathcal{G}$. In particular, assuming the rational injectivity of the Baum–Connes assembly map, one recovers the rational split-injectivity of the comparison map. As coarse groupoids are, in general, not second-countable, one has to provide an additional limiting argument akin to the one used in [5, Section 5.3] to apply the machinery of [19]. This might serve as an alternative proof of the split-injectivity of the rational comparison map for coarse groupoids whose Baum–Connes assembly map is rationally injective. For example, by the results of [23, 26], the class $\mathcal{C}(\mathrm{CE})$ of uniformly locally finite metric spaces admitting a coarse embedding into a Hilbert space satisfies the Baum–Connes conjecture with coefficients; in particular, for these spaces, the Baum-Connes assembly map is rationally injective.
Remark 4.3.
Let $\mathcal{G}$ be a principal, Hausdorff, ample groupoid with compact unit space and let
$\mu\colon K^{\mathrm{top}}(\mathcal{G};C(\mathcal{G}^{(0)}))\to K_{0}(C^{*}_{r}(\mathcal{G}))$
denote the Baum–Connes assembly map for $\mathcal{G}$ with trivial coefficents. It can be shown that there exists a natural group homomorphism
$\alpha^{0}\colon H_{0}(\mathcal{G};\mathbb{Z})\to K^{\mathrm{top}}(\mathcal{G};C(\mathcal{G}^{(0)})),$
such that the 0th comparison map $\alpha_{0}$ factors through $\alpha^{0}$ and the Baum-Connes map. In the case of coarse groupoids, Theorem 4.1 establishes an unconditional low-dimensional injectivity of the Baum-Connes map. We shall return to this perspective in our future work.